# Duoduo CLIP: Efficient 3D Understanding with Multi-View Images

### v1

@@ Abstract
We introduce Duoduo CLIP, a model for 3D representation learning that learns shape encodings from multi-view images instead of point-clouds. The choice of multi-view images allows us to leverage 2D priors from off-the-shelf CLIP models to facilitate fine-tuning with 3D data. Our approach not only shows better generalization compared to existing point cloud methods, but also reduces GPU requirements and training time. In addition, we modify the model with cross-view attention to leverage information across multiple frames of the object which further boosts performance. Compared to the current SOTA point cloud method that requires 480 A100 hours to train 1 billion model parameters we only require 57 A5000 hours and 87 million parameters. Multi-view images also provide more flexibility in use cases compared to point clouds. This includes being able to encode objects with a variable number of images, with better performance when more views are used. This is in contrast to point cloud based methods, where an entire scan or model of an object is required. We showcase this flexibility with object retrieval from images of real-world objects. Our model also achieves better performance in more fine-grained text to shape retrieval, demonstrating better text-and-shape alignment than point cloud based models.
@@ 1 Introduction
Learning representations for 3D shapes is increasingly useful with growing applications in object generation, shape understanding, grounding and robotics. In recent years, there has also been growing interest in 3D representations that are semantically aligned with text. By having a 3D representation aligned to language, these representations can easily be used for text-to-3D shape retrieval [36, 41, 45] and generation [4], or as a basis for 3D feature maps that can be queried with text [16, 25, 19].
Initial work in this domain [4, 36, 41] attempted to train such aligned models on limited data. More recent work [49, 25] leveraged large pre-trained vision-language models (e.g. CLIP [35]) that already align the text and image embeddings. ULIP [49] used frozen CLIP text-image encoders, and trained only the point-cloud encoder but was still limited due to the limited number of 3D objects as well as paired text captions. With larger 3D datasets like Objaverse [7] becoming available in conjunction with automated captioning methods like BLIP-2 [23], recent works [50, 25, 33, 60] trained on such datasets have showcased better performance and stronger generalizability. Despite this, more and more GPU resources are required for training such methods. These recent works [50, 25, 33] also predominantly choose point clouds as their 3D representation. The point clouds are typically limited in resolution and have a domain gap with real-world images, and thus these point cloud based methods can not effectively leverage the 2D priors from CLIP. Uni3D [60] proposed a training regime for incorporating 2D priors with point cloud as input. However, their model requires significant scaling (1B parameters) for the best performance. In this work, we instead choose multi-view images as our shape representation. This not only allows for an easier extension and a more natural way to adapt off-the-shelf 2D representation models. Our Duoduo CLIP also results in more efficient training. As shown in Figure 1(b), our model not only has much stronger generalization on shapes that were not seen during training with a modest model size, but also significantly lowers the amount of compute resources required for training (see Table 1). While some recent work [39, 24] also explore using multi-images with CLIP, their methods do not investigate the use of contrastive learning and only focus on zero/few-shot performance on small-scale datasets.
(b) Accuracy on Objaverse-LVIS with all models trained on Ensembled (no LVIS). Our Duoduo CLIP is fast to train, relatively small while also achieves the best performance.
Figure 1: We propose Duoduo CLIP, a simple model for 3D shape understanding that represents the shape with a multi-view images encoder instead of a point-cloud (PC) encoder. Compared to recent PC-based methods (e.g. OpenShape [25], Uni3D [60]), our model is both accurate and efficient.
@@
OpenShape [25] 1$\times$×A100 (80GB) 300 hr
Uni3D [60] 24$\times$×A100 (40GB) 20 hr
RECON++ [33] 8$\times$×A800 (80GB) 1 day
@@
Using multi-view images also offers greater flexibility in many scenarios. Point cloud based methods require a fairly complete scan of the object, which can be challenging to obtain. In contrast, our model can encode objects from a variable number of frames (1 to M), with more frames capturing more details. Note that we use frames and views interchangeably. This makes our model plug-and-play for any method using image features for encoding 3D objects, enhancing performance with multiple frames. Additionally, previous methods that only train on synthetic objects have limited performance on real-world objects. We further train both our model point cloud based methods on real photos and densely reconstructed objects. Due to artifacts in the point cloud reconstruction pipeline, our model outperforms on real-world objects. Performance is boosted with additional multi-view images where point clouds aren’t available, demonstrating our model’s potential for scaling as more data becomes available.
@@
In summary, our Duoduo CLIP model provides the following: Simple model. We average embeddings across views of an object to obtain 3D representations, while modifying the attention layers to span multiple frames of the shape to enable information flow across frames in the model. Better efficiency. We explore several strategies for efficiently fine-tuning off-the-shelf CLIP models, including training only the attention layers and selectively choosing the number of layers to train without hurting performance. More flexible representation. We showcase benefits of choosing multi-view images as a more flexible 3D representation over point clouds on real-world objects. Furthermore, we also demonstrate better text to shape retrieval with more fine-grained descriptions.
@@ 2 Related Work
There is a large literature of work on 3D representation learning from 3D data either using supervised training [31, 34] or using self-supervised learning (with reconstruction loss [5], contrastive learning [48], masked prediction [54], and others). Here we focus on work that learns aligned representation spaces between text and either 2D or 3D.
@@
2D-text joint representation. Modeling multiple modalities through a shared representation has been explored through early works [10, 11, 21, 47] by aligning images and text embeddings. With the introduction of the InfoNCE loss [29], ConVIRT [57] proposed using contrastive learning with InfoNCE style loss to learn joint text and image embeddings. This was popularized by follow-up work such as CLIP [35] and ALIGN [20], which demonstrated that by scaling up training on a large-scale dataset, the learned embeddings can be useful for many downstream tasks such as classification, retrieval, grounding and generation. Followup work [20, 51, 28, 55, 52] has since been proposed to improve upon the formulation or training. There have also been attempts such as MERU [8] that use a hyperbolic embedding space to better model the hierarchical relations between text and images.
@@
3D-text joint representation. Text2Shape [4] was one of the first works that introduced a captioned text and 3D dataset and used metric learning to learn an aligned space between text and 3D shapes. Follow-up work investigated how to improve the aligned representation for text-to-shape retrieval with triplet loss [41] and contrastive losses [36, 49, 45]. Prior work also investigated the use of CLIP to align the embeddings [43, 36, 49]. ULIP [49] used frozen CLIP text-image encoders, and trained only the point-cloud encoder. However, the generalizability of these works were limited by the available dataset size at the time. With the introduction of larger scale 3D datasets such as Objaverse [7], and the use of LVLMs for captioning 3D datasets [25, 27], more recent work [25, 50] were able to train with larger datasets of paired text and 3D shapes. OpenShape [25] and ULIP-2 [50] were among the first to use such datasets to train a point cloud encoder aligned with text using contrastive learning. TAMM [58] leverages additional adapters to better align with the pre-trained CLIP models. Uni3D [60] exploits 2D priors within pre-trained large-scale vision models and maps point cloud patches to image patches within pre-trained ViTs to enable the scaling of even larger models. VIT-LENS [22] follows a similar pipeline by training a point cloud embedding layer and perceiver model to map to a frozen CLIP model. RECON++ [33] similar to RECON [32] employs both contrastive and MAE [15] reconstruction losses with an additional multi-view image matching cost. These recent works typically rely on frozen text and image encoders, and all used point-cloud encoders. Our work shows that taking embeddings of multi-view images is more efficient and works better with real-world images.
@@
Multi-view images for text-3D representations. Multi-view images have long been used to represent 3D shapes [2, 37, 40, 46, 13], with recent work [39, 24] using multi-view images for shape representation by aggregating CLIP features. While these works also use CLIP with multi-view images, our focus is whether using a trainable multi-view shape encoder as a drop-in-replacement for the point-cloud encoder in contrastive settings such as OpenShape will result in better shape-representation for downstream tasks. These works do not focus on improved representation learning, but on how to use prompting to select better views or view aggregation and improve shape classification. There is also a line of work [56, 61, 17] that uses multi-view depth maps from point cloud to adapt CLIP models for 3D. However, the focus of these methods is still on point clouds and the models show limited performance on large-scale datasets [50, 25].
@@ 3 Method
Figure 2: Our multi-view encoder takes a variable number $[1,M]$ of images as input and outputs a single shape embedding. The first few layers of the ViT model are frozen (blue) and attention operates on the individual frames in parallel. For the latter layers, the attention layers are trainable (orange) and modified to attend over all $M$ views. The embeddings for each frame are averaged to get the final embedding. Note that only major components of the model are depicted.
Our Duoduo CLIP learns a shape representation by using contrastive learning to encode multi-view images into a pre-aligned text-image space. Our contrastive learning framework is similar to previous works [49, 50, 25, 58, 60], with a shape encoder $E^{S}$ and frozen image $E^{I}$ and text $E^{T}$ encoders from a pre-trained CLIP model. The key difference of our model is that we replace the point cloud based shape encoder with one that uses multi-view images. Figure 1(a) shows a comparison of our model to prior work, and Figure 2 shows the details of our multi-view shape encoder.
To perform training, the dataset is organized into triplets of $(S_{i},I_{i},T_{i})$, where $S$ is the shape representation, $I$ are rendered images of the 3D object and $T$ is the corresponding text description. Embeddings are calculated for each modality $f^{S}_{i}=E^{S}(S_{i})$, $f^{I}_{i}=E^{I}(I_{i})$ and $f^{T}_{i}=E^{T}(T_{i})$. The asymmetric contrastive loss [57] between any two modalities is:
@@
For both $E^{S}$ and $E^{I}$ we average the extracted embeddings for all $m$ views to produce a single latent representation for an object: $f^{S}_{i}=E^{S}(S_{i})=\frac{1}{m}\Sigma_{j=1}^{m}\text{ViT}(S_{i})[j]$, where ViT [9] is the backbone encoder. Here we abuse the notation a bit, since our method allows interaction between views in the model, the index operation gets the corresponding embedding from the $j$-th frame’s [CLS] token. We initialize $E_{S}$ initially from $E^{I}$. During training, we randomly sample $1$ to $M$ views for each step so that the model is robust to a variable number of views.
Multi-view attention. We incorporate information flow between views by modifying the self-attention layers to attend over tokens across multiple views similar to recent multi-view aware methods like MVDream [38]. The original self-attention formulation [44] is $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}}V)$, where $Q,K,V\in R^{L\times k}$, and $L$ is the token length for a single image after patchify operations and $k$ is the channel dimension. For cross-view attention we aggregate token features across $m$ multi-views such that $Q^{\text{MV}},K^{\text{MV}},V^{\text{MV}}\in R^{(L\times m)\times k}$ (MV stands for multi-view) and the $\operatorname{Attention}(Q^{\text{MV}},K^{\text{MV}},V^{\text{MV}})$ can be calculated without changing the parameters of the original model. After attention, the views are separated and processed in parallel in other parts of the model. This mechanism helps the model reason between different views of the object and aggregate the most important information across all tokens.
Trainable layers. To reduce the memory needed for fine-tuning, we only train the self-attention layers. We further cut down on memory usage by reducing the number of trainable layers. Since the non-trainable self-attention layers have only been trained to see patches of a single image, we only enable the multi-view attention layers for the unfrozen attention layers (see Figure 2).
Figure 3: Left. We compare the Top 1 accuracies on Objaverse-LVIS for our model (Duoduo CLIP) and two pretrained CLIP models (zero shot) for different number of frames. B-32 is the model we initialize from and train with and BigG-14 is the model used by OpenShape. The dashed lines show the difference between using 1 and 12 frames. Right. We show the Top1, Top3 and Top5 accuracies for our model for 1 to 12 frames (note the y-axis starts at 40). Gray dashed lines shows the inference time(s) for 1 iteration and batch size 200 at different number of frames.

### v3

@@ Abstract
We introduce Duoduo CLIP, a model for 3D representation learning that learns shape encodings from multi-view images instead of point clouds. The choice of multi-view images allows us to leverage 2D priors from off-the-shelf CLIP models to facilitate fine-tuning with 3D data. Our approach not only shows better generalization compared to existing point cloud methods, but also reduces GPU requirements and training time. In addition, the model is modified with cross-view attention to leverage information across multiple frames of the object which further boosts performance. Notably, our model is permutation invariant to the order of multi-view images while being pose-free. Compared to the current SOTA point cloud method that requires 480 A100 hours to train 1 billion model parameters we only require 57 A5000 hours and 87 million parameters. Multi-view images also provide more flexibility including being able to encode objects with a variable number of images, and performance scales when more views are used. In contrast, point cloud based methods require an entire scan or model of the object. We showcase this flexibility with benchmarks from images of real-world objects. Our model also achieves better performance in more fine-grained text to shape retrieval, demonstrating better text-and-shape alignment than point cloud based models.
@@ 1 Introduction
Learning representations for 3D shapes is useful for many applications in object generation, shape understanding, grounding and robotics. Recently, there is also growing interest in aligning 3D representations to text, to enable text-to-3D shape retrieval (Ruan et al., 2024, Tang et al., 2023, Wang et al., 2023) and generation (Chen et al., 2019), or as a basis for 3D feature maps that can be queried with text (Huang et al., 2023a, Liu et al., 2023a, Jatavallabhula et al., 2023).
@@
OpenShape Liu et al. (2023a) 1$\times$×A100 (80GB) 300 hr
Uni3D Zhou et al. (2024) 24$\times$×A100 (40GB) 20 hr
RECON++ Qi et al. (2024) 8$\times$×A800 (80GB) 1 day
@@
Early work in this domain attempted to train such aligned models on limited data (Chen et al., 2019, Ruan et al., 2024, Tang et al., 2023). More recent work (Xue et al., 2023, Liu et al., 2023a) leveraged large pre-trained vision-language models (e.g. CLIP (Radford et al., 2021)). Larger 3D datasets like Objaverse (Deitke et al., 2023) enabled training of methods that can handle more diverse object categories (Xue et al., 2024, Liu et al., 2023a, Qi et al., 2024, Zhou et al., 2024). However, more and more GPU resources are required for training such methods. These works (Xue et al., 2024, Liu et al., 2023a, Qi et al., 2024) predominantly rely on point clouds as their 3D representation. Point clouds are limited in resolution and have a domain gap with real-world images, and thus these point cloud-based methods can not effectively leverage 2D priors from CLIP. Uni3D (Zhou et al., 2024) proposed a training regime to incorporate 2D priors with point clouds as input. However, their model requires significant scaling (1B parameters) for the best performance. In this work, we show that our Duoduo CLIP approach which uses multi-view images as the shape representation allows more easily adapting off-the-shelf 2D representation models while enabling more efficient training. As shown in Fig. 1(b), our model not only has much stronger generalization on shapes that were not seen during training, but also significantly lowers the compute resources required for training (see Tab. 1).
(b) Accuracy on Objaverse-LVIS with all models trained on Ensembled (no LVIS). Our Duoduo CLIP is fast to train, relatively small while also achieving the best performance.
Figure 1: Duoduo CLIP is a model for 3D shape understanding that represents the shape with a multi-view images encoder instead of a point-cloud (PC) encoder. This makes our model more accurate and efficient than PC methods (OpenShape Liu et al. (2023a), Uni3D Zhou et al. (2024)).
Figure 2: We blend concepts from two distinct objects by identifying the Objaverse shape with the embedding that maximizes similarity to both objects simultaneously. This shows that our model has learned a latent space capable of combining geometric features from the mixed objects or capturing their semantic attributes. For more details see Sec. A.4.2.
@@
Multi-view images also offer greater flexibility. For example, in real-world datasets like ScanNet (Dai et al., 2017) or ScanObjectNN (Uy et al., 2019) a 3D mesh is first reconstructed from densely captured RGBD data from which 3D point clouds are sampled. This process introduces errors and artifacts, in contrast to sparse multi-view images which can be directly sub-sampled from the RGB video stream. Our model takes images directly as input, making it plug-and-play for any method using image features for encoding 3D objects, which is particularly important for real-time robotics applications with sparse views (Wang et al., 2024, Gadre et al., 2023, Garg et al., 2024, Shridhar et al., 2022). Additionally, previous methods (Liu et al., 2023a) that only train on synthetic objects have limited performance on real-world objects. Our method can be trained on both real photos and reconstructed objects and outperforms prior work on real-world objects. Performance can further be boosted with additional multi-view images when point clouds are not available, demonstrating our model’s potential for scaling as more data becomes available.
@@
In summary, our Duoduo CLIP model provides the following: Simpler model. We average embeddings across views of an object to obtain 3D representations, while modifying the attention layers to span multiple frames of the shape to enable information flow across frames in the model. Better efficiency. We explore several strategies for efficiently fine-tuning off-the-shelf CLIP models, including training only the attention layers and selectively choosing the number of layers to train without hurting performance. More flexible representation. We show the benefits of multi-view images as a more flexible 3D representation over point clouds on real-world objects. Furthermore, we demonstrate better text to shape retrieval with more fine-grained descriptions, and that by maximizing similarity to two different shape embeddings, our learned embedding space can also be used to blend concepts (see Fig. 2).
@@ 2 Related Work
There is much work on 3D representation learning using supervised training (Qi et al., 2017, Qian et al., 2022) or using self-supervised learning (with reconstruction loss (Chen & Zhang, 2019), contrastive learning (Xie et al., 2020), masked prediction (Yu et al., 2022b), and others). Here we focus on work that learns aligned representation spaces between text and either 2D or 3D.
@@
2D-text joint representation. Modeling multiple modalities through a shared representation has been explored through early works by aligning images and text embeddings (Faghri et al., 2018, Frome et al., 2013, Kiros et al., 2014, Weston et al., 2011). With the introduction of the InfoNCE loss (Oord et al., 2018), ConVIRT (Zhang et al., 2022b) proposed using contrastive learning with InfoNCE style loss to learn joint text and image embeddings. This was popularized by follow-up work such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), which demonstrated that by scaling up training on a large-scale dataset, the learned embeddings can be useful for many downstream tasks such as classification, retrieval, grounding and generation. Followup work (Jia et al., 2021, Yang et al., 2022, Mu et al., 2022, Zhai et al., 2022, Yu et al., 2022a) has since improved upon the formulation or training. There have also been attempts such as MERU (Desai et al., 2023) that use a hyperbolic embedding space to better model the hierarchical relations between text and images.
@@
3D-text joint representation. Text2Shape (Chen et al., 2019) was one of the first works that introduced a captioned text and 3D dataset and used metric learning to learn an aligned space between text and 3D shapes. Follow-up work investigated improvements to the aligned representation for text-to-shape retrieval with triplet loss (Tang et al., 2023) and contrastive losses (Ruan et al., 2024, Xue et al., 2023, Wang et al., 2023). Some prior work used of CLIP to align the embeddings (Thomason et al., 2022, Ruan et al., 2024, Xue et al., 2023). ULIP (Xue et al., 2023) used frozen CLIP text-image encoders, and trained only the point-cloud encoder. However, generalization ability was limited by the available dataset size at the time. With the introduction of larger scale 3D datasets such as Objaverse (Deitke et al., 2023), and the use of LVLMs for captioning 3D datasets (Liu et al., 2023a, Luo et al., 2024), more recent work (Liu et al., 2023a, Xue et al., 2024) trained with larger datasets of paired text and 3D shapes. OpenShape (Liu et al., 2023a) and ULIP-2 (Xue et al., 2024) were among the first to use such datasets to train a point cloud encoder aligned with text using contrastive learning. TAMM (Zhang et al., 2024) leverages additional adapters to better align with the pre-trained CLIP models. Uni3D (Zhou et al., 2024) exploits 2D priors within pre-trained large-scale vision models and maps point cloud patches to image patches within pre-trained ViTs to enable the scaling of even larger models. VIT-LENS (Lei et al., 2024) follows a similar pipeline by training a point cloud embedding layer and perceiver model to map to a frozen CLIP model. RECON++ (Qi et al., 2024) similar to RECON (Qi et al., 2023) employs both contrastive and MAE (He et al., 2022) reconstruction losses with an additional multi-view image matching cost. These recent works typically rely on frozen text and image encoders, and all used point-cloud encoders. Our work shows that taking embeddings of multi-view images is more efficient and works better with real-world images. Note that Gao et al. (2024) combines point clouds with optional multi-view embeddings from a frozen CLIP encoder, which lacks inter-view context, and does not support arbitrary views or poses, resulting in lower performance at the same number of views while also requiring point cloud.
@@
Multi-view images for text-3D representations. Multi-view images have long been used to represent 3D shapes (Bradski & Grossberg, 1994, Shen et al., 2003, Su et al., 2015, Wei et al., 2020, Hamdi et al., 2021), with recent work (Song et al., 2023, Lin et al., 2024) using multi-view images for shape representation by aggregating CLIP features. While these works also use CLIP with multi-view images, our focus is whether using a trainable multi-view shape encoder as a drop-in-replacement for the point-cloud encoder in contrastive settings such as OpenShape will result in better shape representations for downstream tasks. These works do not focus on improved representation learning, but on how to use prompting to select better views or view aggregation and improve shape classification. There is also a line of work (Zhang et al., 2022a, Zhu et al., 2023, Huang et al., 2023b) that uses multi-view depth maps from point cloud to adapt CLIP models for 3D. However, the focus of these methods is still on point clouds and the models show limited performance on large-scale datasets (Xue et al., 2024, Liu et al., 2023a).
@@ 3 Method
Figure 3: Our multi-view encoder takes a variable number $[1,M]$ of images as input and outputs a single-shape embedding. The first few layers of the ViT model are frozen (blue) and attention operates on the individual frames in parallel. For the latter layers, the attention layers are trainable (orange) and modified to attend over all $M$ views. The embeddings for each frame are averaged to get the final embedding. Note that only major components of the model are depicted.
Our Duoduo CLIP learns a shape representation by using contrastive learning to encode multi-view images into a pre-aligned text-image space. Our contrastive learning framework is similar to previous works (Xue et al., 2023; 2024, Liu et al., 2023a, Zhang et al., 2024, Zhou et al., 2024), with a shape encoder $E^{S}$ and frozen image $E^{I}$ and text $E^{T}$ encoders from a pre-trained CLIP model. The key difference is that we replace the point cloud shape encoder with one that uses multi-view images. Fig. 1(a) illustrates this difference, while Fig. 3 shows the details of our multi-view shape encoder.
To perform training, the dataset is organized into triplets of $(S_{i},I_{i},T_{i})$, where $S$ is the shape representation, $I$ are rendered images of the 3D object and $T$ is the corresponding text description. Embeddings are calculated for each modality $f^{S}_{i}=E^{S}(S_{i})$, $f^{I}_{i}=E^{I}(I_{i})$ and $f^{T}_{i}=E^{T}(T_{i})$. The asymmetric contrastive loss (Zhang et al., 2022b) between any two modalities is:
@@
Two important properties of our shape encoder for multi-view images include being able to handle arbitrary number of views as well as being permutation invariant to ensure the flexibility for the input images. Encoding of any number of frames with the model into a single embedding as well as a permutation invariant operation between multi-views are discussed in more detail below. Note that we also do not make assumptions about the poses of the input image, making our method pose-free.
@@
For both $E^{S}$ and $E^{I}$ we average the extracted embeddings for all $m$ views to produce a single latent representation for an object: $f^{S}_{i}=E^{S}(S_{i})=\frac{1}{m}\Sigma_{j=1}^{m}\text{ViT}(S_{i})[j]$, where ViT (Dosovitskiy et al., 2021) is the backbone encoder. Here we abuse the notation a bit, since our method allows interaction between views in the model, the index operation gets the corresponding embedding from the $j$-th frame’s [CLS] token. We initialize $E_{S}$ initially from $E^{I}$. During training, we randomly sample $1$ to $M$ views for each step so that the model is robust to a variable number of views.
Figure 4: Attention maps are extracted from the first trainable MVA layer, capturing token-wise attention across all M views. A specific token (e.g., the plate) is queried to retrieve the corresponding row from the attention map, and its magnitudes are mapped onto the original images for visualization. The model has learned geometric correspondences of 3D shapes across different multi-views.
Multi-view attention. We incorporate information flow between views by modifying the self-attention layers to attend over tokens across multiple views similar to recent multi-view aware methods like MVDream (Shi et al., 2023). The original self-attention (Vaswani et al., 2017) is given by $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$, where $Q,K,V\in R^{L\times k}$, $L$ is the token length for a single image after patchify operations, and $k$ is the channel dimension. For cross-view attention we aggregate token features across $m$ multi-views (MV) so that $Q^{\text{MV}},K^{\text{MV}},V^{\text{MV}}\in R^{(L\times m)\times k}$ and the $\operatorname{Attention}(Q^{\text{MV}},K^{\text{MV}},V^{\text{MV}})$ can be calculated without changing the parameters of the original model. After attention, the views are separated and processed in parallel in other parts of the model. This mechanism helps the model reason between different views of the object and aggregate the most important information across all tokens. In Fig. 4, we visualize the attention across views corresponding to a selected token (see Sec. A.5.1 for details). From the figure, we see the multi-view attention layers learn to capture correspondence between object parts from different views.
Trainable layers. To reduce the memory needed for fine-tuning, we only train a portion of the self-attention layers while keeping the MLP layers after attention fixed. Since the non-trainable self-attention layers have only been trained to see patches of a single image, we only enable the multi-view attention layers for the unfrozen attention layers (see Fig. 3). Additionally, we find the choice of freezing the MLPs, helps in preventing overfitting and makes the model generalize better on unseen datasets (see Sec. A.2.2).